# references.bib A comparison of methods for estimating the average treatment effect on the treated for externally controlled trials

### v1

@@ Abstract
While randomized trials may be the gold standard for evaluating the effectiveness of the treatment intervention, in some special circumstances, single-arm clinical trials utilizing external control may be considered. The causal treatment effect of interest for single-arm studies is usually the average treatment effect on the treated (ATT) rather than the average treatment effect (ATE). Although methods have been developed to estimate the ATT, the selection and use of these methods require a thorough comparison and in-depth understanding of the advantages and disadvantages of these methods. In this study, we conducted simulations under different identifiability assumptions to compare the performance metrics (e.g., bias, standard deviation (SD), mean squared error (MSE), type I error rate) for a variety of methods, including the regression model, propensity score matching, Mahalanobis distance matching, coarsened exact matching, inverse probability weighting, augmented inverse probability weighting (AIPW), AIPW with SuperLearner, and targeted maximum likelihood estimator (TMLE) with SuperLearner.
@@
According to our simulation results under most scenarios we examined, propensity score matching methods perform best overall in type I error control. However, they in general have worse performance in the estimation accuracy compared to doubly robust methods given that the identifiability assumptions are not severely violated.
Keywords single-arm studies external control real-world data causal inference average treatment effect on the treated
@@ 1 Introduction
Randomized clinical trials (RCTs) are usually considered the gold standard for demonstrating the efficacy and safety of an intervention. However, for rare diseases it is difficult to recruit a large number of patients, and for severe diseases it is unethical to assign patients to placebo. Single-arm studies utilizing an external real-world data (RWD) control could be an appealing option. In addition to the feasibility and ethical advantages, single-arm trials with external RWD control may save time and money and as a result have been receiving increasing attention.
To draw inferences about the effect of a treatment, a target causal effect needs to be defined. For the single-arm studies with external RWD control, the causal treatment effect of interest could be the average treatment effect on the treated (ATT) (i.e., the average additive treatment effect for the population receiving the treatment compared to what would have happened if they had not received treatment) if the treatment effect among the enrolled population is of more interest [heckman1985alternative, heckman1997matching]. This is particularly the case if there is heterogeneity between patients participating and those not participating in the trial due to eligibility criteria. For example, if the experimental drug is a second-line treatment, the study drug’s effect cannot be considered for naive patients who are not eligible to participate in the trial.
@@
This paper is organized as follows. The introduction section provides background information on single-arm trials with external RWD control and presents the motivation behind this study. The Methods section revisits the definitions of ATT and identifiability assumptions and provides a description of the models and simulation settings utilized in this study. The Results section presents the findings from various scenarios. In the Discussion section, we analyze the experimental results and provide recommendations for planning single-arm trials with external control. Finally, the Conclusion section summarizes the study’s key findings, interpretation, and implications.
@@ 2.1 Average treatment effect on the treated
In a real-world study, a causal effect of a treatment can be defined in terms of the comparison between potential outcomes (also known as the counterfactual outcomes) of a population who would receive the treatment and potential outcomes of the same population who would not receive the treatment [hernan2020causal]. The comparison is drawn from a population level instead of an individual level because only one of the potential outcomes can be observed for each patient [abadie2005semiparametric]. In a nonrandomized, single-arm clinical trial with external RWD control, the population of interest is usually only the patients receiving the treatment [fang2020statistical]. Correspondingly, the causal effect of interest is the ATT.
For patient with index , let denote the patient’s treatment assignment with a value of 1 for treated and 0 for not treated, vector of the patient’s covariates, the patient’s observed outcome, and the patient’s potential outcomes corresponding to and , respectively. Then the ATT on the difference scale , which is the causal effect focused on in this study, can be formally defined as
@@ 2.2 Identifiability assumptions
The positivity assumption implies that all subjects have a non-zero probability of receiving (or not receiving) the treatment. Positivity violations can be divided into two categories: 1) theoretical violation that the probability of being treated is 0 or 1, e.g., certain patients are restricted from receiving the treatment or control, and 2) near or practical violation in the sampling due to high/low exposure prevalence or small sample size [leger2022causal]. One method for assessing the positivity assumption is via the propensity score (PS), which is the probability of receiving the treatment conditional on covariate values, i.e., [zhu2021core]. In particular, we can study the distribution of the PS for the study population in which subjects with PS values near 0 and 1 are more likely to violate the positivity assumption (Figure 1). If many PS values are distributed at 0 or 1, then the positivity assumption cannot hold; on the other hand, if no PS values are distributed around 0 or 1, the positivity assumption can be considered reasonable.
@@ 2.3.1 Linear regression model
In general, a regression model is not a sound model for estimating ATT. However, if the true outcome model can be properly fitted by a regression model, the regression model may estimate ATT well. In particular, if 1) the identifiability assumptions are met and 2) the outcome model is correctly specified, the ATT can be estimated through the regression model. However, in practice, it is challenging to specify the correct outcome model. In this study, we use the linear regression that includes the specified covariates and the treatment variable without interaction terms as a simple reference model. The ATT is estimated through the coefficient of the treatment variable .
@@ 2.3.3 Coarsened exact matching
In practice, exact matching of covariates may not be feasible when there are continuous covariates or when there are categorical covariates with many levels. One approach to address this issue is the CEM method, which is a form of exact matching based on the subclasses of covariates [iacus2011multivariate]. It involves two steps: 1) coarsen each covariate into several “bins,” i.e., subclasses, and 2) perform exact matching on those bins. The unmatched units are discarded.
For CEM, the choice of the number of bins is crucial to model performance. Domain knowledge is required to determine the appropriate number of bin for each variable [iacus2011multivariate]. In this study, we arbitrarily choose 2 (denoted by CEM2), a smaller number of bin, and 5 (denoted by CEM5), a reasonably large one, for all covariates to study the performance of the CEM model with different numbers of bins. For each number of bins, the boundaries of the bins were set to be evenly spaced between the maximum and minimum values of the covariate [ho2018package].
@@ 2.3.8 Targeted maximum likelihood estimator
The application of TMLE for estimating ATT involves the following steps [van2011targeted]:
@@
Using the empirical distribution of on the treated population, , obtain an initial ATT estimate
@@
Starting with , , and , iteratively update the estimate of , , and until the following efficient score equation is solved
where
and . Then is the TMLE of .
In this study, we implemented TMLE by using the SuperLearner for the estimation of both propensity scores and outcomes [gruber2006package]. The model is denoted as TMLE_SL. In addition, we implemented an model, denoted as AIPW_SL, on the basis on the AIPW method and SL’s estimation of the propensity scores and outcomes.
@@ 2.4 Simulation
In Scenarios 1 and 2, we assumed three independent covariates that followed the standard normal distribution. was a confounder that affected both the treatment selection and the outcome, and and affected the treatment selection and the outcome, respectively. In Scenario 3, an additional confounder was assumed that also followed the standard normal distribution and was independent of .
@@
Setting 3: Outcome model with interaction terms between the treatment and covariates.
@@
Estimated the ATT by using the methods described in Section 2.3. For each method, we included only covariates for estimation. We used the logistic regression based on to estimate the PS for PSM, PSM_1:2, IPW, and AIPW. And we used the linear regression based on to estimate the outcomes for AIPW.
@@
The type I error rate, i.e., the average rate of rejecting the null hypothesis of by making in the outcome models described in Table 2.
@@ 3.1 Scenario 1: no violation of the identifiability assumptions
The distribution of the PS values for Scenario 1 assuming no violation of the positivity or the unconfoundedness is shown in Figure 4. As shown in the figure, the distributions are bounded away from 0 and 1, and thus the positivity assumption can be considered reasonable. The bias, empirical SD (hereafter referred to as SD), average theoretical SD, MSE, and type I error rate for the 10 methods in Scenario 1 under three different settings are shown in Figure 5, Figure 6, and Figure 7, respectively.
@@
In Setting 3 with the interaction between the treatment and covariates in the true outcome model, most methods yield very small biases. In terms of SD, the matching-based methods in general have relatively larger SDs than the other methods. The MSE pattern is similar to the SD pattern. Under the null treatment effect, the outcome models for Settings 1 and 3 are the same, therefore, the findings regarding the type I error are also the same for these two settings.
@@ 3.2 Scenario 2: violation of the positivity assumption
The distribution of the PS values for Scenario 2 assuming violation of the positivity or the unconfoundedness is shown in Figure 4. As shown in the figure, a large proportion of subjects’ PS values are distributed at or close to 0 or 1 and thus the positivity assumption is considered violated. The bias, empirical SD, average theoretical SD, MSE, and type I error rates for the 10 methods in Scenario 2 under three different settings are shown in Figure 8, Figure 9, and Figure 10, respectively.
@@
In Setting 1 with linear effect in the true outcome models, CEM2 and IPW in general have noticeably larger biases compared to other methods. The increased bias of CEM2 in Scenario 2 compared to Scenario 1 is attributed to the impact of violating the positivity assumption. Under Scenario 2, many subjects cannot be matched exactly but were still matched when using two bins, thus causing a larger bias. The IPW also produces a relatively larger bias under Scenario 2. The SD pattern in Scenario 2 differs from that in Scenario 1 in the following ways: 1) the SDs of IPW, AIPW, and TMLE_SL have a noticeable increase, and 2) the average theoretical SD of TMLE_SL is much smaller than the empirical SD. In terms of MSE, LR and AIPW_SL are the best performing models. Under the null treatment effect, CEM2, IPW, and TMLE_SL experienced remarkable type I error inflation. According to Figure 8, the type I error inflation of the CEM2 or IPW is mainly due to the large bias, while the type I error inflation of the TMLE is mainly due to its underestimation of the SD. In contrast, PSM and PSM_1:2 perform best in terms of type I error control.
@@
In Setting 2 with nonlinear effect in the true outcome models, in addition to the CEM2 and IPW, which are already substantially biased in Setting 1, LR and AIPW are also largely biased. The bias of CEM5 is relatively large when the prevalence rate of treatment is low and relatively small when the prevalence rate of treatment is high. The biases of PSM and PSM_1:2 are the smallest. The SDs of most models in Setting 2 increase compared to Setting 1, especially for IPW and LR. The AIPW_SL has a relatively small bias and the smallest SD, resulting the smallest MSE. Under the null treatment effect, LR, CEM2, IPW, AIPW, and CEM5 show severe type I error inflation, mainly due to their large biases (Figure 9); TMLE_SL also shows severe type I error inflation, mainly due to its underestimation of the SD (Figure 9). AIPW_SL and MDM have relatively mild type I error inflation, while PSM and PSM_1:2 have the best type I error control.
@@
In Setting 3 with the interaction between the treatment and covariates in the true outcome model, the IPW, LR, and matching-based methods in general have relatively large biases (Figure 10). The doubly robust methods (i.e., AIPW, AIPW_SL, and TMLE_SL) in general have relatively small biases. Among the doubly robust methods, AIPW and TMLE_SL have relatively large SDs and hence relatively large MSEs. The MSE of AIPW_SL is the smallest, due to its relatively small bias and SD. The findings for Setting 3 regarding the type I error inflation are the same as those for Setting 1 (Figure 10).
@@ 3.3 Scenario 3: violation of the unconfoundedness assumption
For all methods and three settings of the outcome model, violation of the unconfoundedness assumption in Scenario 3 leads to a certain amount of bias (or a larger bias if the method is already biased) compared to Scenario 1 where this assumption is not violated. The biases induced by violation of unconfoundedness assumption are all positive under our simulation settings. We conducted post-hoc analyses by varying the coefficient values of the unmeasured confounder in the three settings and found that the sign and magnitude of the induced bias are mainly determined by the coefficient value of the unmeasured confounder.
The pattern of SD is similar between Scenario 1 and Scenario 3. The MSEs observed in Scenario 3 are greater than those observed in Scenario 1. Under the null hypothesis, the type I error rates in Scenario 3 generally increase compared to Scenario 1 due to the larger bias.
@@ 4.1.1 Matching-Based Methods
Among matching-based methods, our simulation results show that the PSM model in general results in a small bias when the positivity assumption is held. When the positivity assumption is violated and when there is interaction between the treatment variable and covariates (i.e., Scenario 2 Setting 3), a mild bias can be induced. This is because under the violation of the positivity assumption, the integration of in equation (2) is carried out in the overlapped range of the PS distributions for the two groups and hence the causal effect of interest has been deviated from the ATT. Similar issues occur in the other matching-based methods that we studied. Of note, the bias of the PSM can still be small in cases with positivity violation but without interaction between the treatment variable and covariates, e.g., Scenario 2 Settings 1 or 2, because in equation (2), is a constant, and its integration over is the same as its integration in the overlapped range of the PS distribution.

### v2

@@ Abstract
While randomized trials may be the gold standard for evaluating the effectiveness of the treatment intervention, in some special circumstances, single-arm clinical trials utilizing external control may be considered. The causal treatment effect of interest for single-arm trials is usually the average treatment effect on the treated (ATT) rather than the average treatment effect (ATE). Although methods have been developed to estimate the ATT, the selection and use of these methods require a thorough comparison and in-depth understanding of the advantages and disadvantages of these methods. In this study, we conducted simulations under different identifiability assumptions to compare the performance metrics (e.g., bias, standard deviation (SD), mean squared error (MSE), type I error rate) for a variety of methods, including the regression model, propensity score matching (PSM), Mahalanobis distance matching (MDM), coarsened exact matching, inverse probability weighting, augmented inverse probability weighting (AIPW), AIPW with SuperLearner, and targeted maximum likelihood estimator (TMLE) with SuperLearner.
@@
According to the simulation results, under most scenarios we examined, PSM and MDM methods perform best overall in terms of type I error control. However, they in general have worse performance in the estimation accuracy compared to doubly robust methods given that the identifiability assumptions are not severely violated.
Keywords Single-arm trials External control Real-world data Causal inference Average treatment effect on the treated
@@ 1 Introduction
Randomized clinical trials (RCTs) are usually considered the gold standard for demonstrating the efficacy and safety of an intervention. However, for rare diseases it is difficult to recruit a large number of patients, and for severe diseases it is unethical to assign patients to placebo. Single-arm trials utilizing an external real-world data (RWD) control could be an appealing option. In addition to the feasibility and ethical advantages, single-arm trials with external RWD controls may save time and money and as a result have been receiving increasing attention.
To draw inferences about the effect of a treatment, a target causal effect needs to be defined. For the single-arm trials with external RWD controls, the causal treatment effect of interest could be the average treatment effect on the treated (ATT) (i.e., the average treatment effect for the population receiving the treatment compared to what would have happened if they had not received treatment) if the treatment effect among the enrolled population is of more interest [heckman1985alternative, heckman1997matching]. This is particularly the case if there is heterogeneity between patients participating and those not participating in the trial due to eligibility criteria. For example, if the experimental drug is a second-line treatment, the study drug’s effect cannot be considered for naive patients who are not eligible to participate in the trial.
@@
This paper is organized as follows. The introduction section provides background information on single-arm trials with external RWD controls and presents the motivation behind this study. The Methods section revisits the definitions of ATT and identifiability assumptions and provides a description of the models and simulation settings utilized in this study. The Results section presents the findings from various scenarios. In the Discussion section, we analyze the experimental results and provide recommendations for planning single-arm trials with external controls. Finally, the Conclusion section summarizes the study’s key findings, interpretation, and implications.
@@ 2.1 Average treatment effect on the treated
In a real-world study, a causal effect of a treatment can be defined in terms of the comparison between potential outcomes (also known as the counterfactual outcomes) of a population who would receive the treatment and potential outcomes of the same population who would not receive the treatment [hernan2020causal]. The comparison is drawn from a population level instead of an individual level because only one of the potential outcomes can be observed for each patient [abadie2005semiparametric]. In a nonrandomized, single-arm clinical trial with an external RWD control, the population of interest is usually only the patients receiving the treatment [fang2020statistical]. Correspondingly, the causal effect of interest is the ATT.
For patient with index , let denote the patient’s treatment assignment with a value of 1 for treated and 0 for not treated, vector of the patient’s covariates, the patient’s observed outcome, and the patient’s potential outcomes corresponding to and , respectively. Then the ATT of our interest in terms of the difference scale rather than others (e.g., ratio), can be formally defined as
@@ 2.2 Identifiability assumptions
The positivity assumption implies that all subjects have a non-zero probability of receiving (or not receiving) the treatment. Positivity violations can be divided into two categories: 1) theoretical violation that the probability of being treated is 0 or 1, e.g., certain patients are restricted from receiving the treatment or control, and 2) near or practical violation in the sampling due to high/low exposure prevalence or small sample size [leger2022causal]. One method for assessing the positivity assumption is via the propensity score (PS), which is the probability of receiving the treatment conditional on covariate values, i.e., [zhu2021core]. In particular, we can study the distribution of the PS for the study population and explore whether there are subjects with PS values near 0 or 1 (Figure 1). If many PS values are distributed around 0 or 1, then the positivity assumption should not hold; on the other hand, if no PS values are distributed around 0 or 1, the positivity assumption can be considered reasonable.
@@ 2.3.1 Linear regression model
In general, a regression model is not a sound model for estimating ATT. However, if the true outcome model can be properly fitted by a regression model, the regression model may estimate ATT well. In particular, if 1) the identifiability assumptions are met and 2) the outcome model is correctly specified, the ATT can be estimated through the regression model. However, in practice, it is challenging to specify the correct outcome model. In this study, we use the linear regression (LR) that includes the specified covariates and the treatment variable without interaction terms as a simple reference model. The ATT is estimated through the coefficient of the treatment variable .
@@ 2.3.3 Coarsened exact matching
In practice, exact matching of covariates may not be feasible when there are continuous covariates or when there are categorical covariates with many levels. One approach to address this issue is the CEM method, which is a form of exact matching based on the subclasses of covariates [iacus2011multivariate]. It involves two steps: 1) coarsening each covariate into several “bins,” i.e., subclasses, and 2) performing exact matching on those bins. The unmatched units are discarded.
For CEM, the choice of the number of bins is crucial to model performance. Domain knowledge is required to determine the appropriate number of bin for each variable [iacus2011multivariate]. In this study, we arbitrarily choose 2 (denoted by CEM2), a smaller number of bin, and 5 (denoted by CEM5), a reasonably larger one, for all covariates to study the performance of the CEM model with different numbers of bins. For each number of bins, the boundaries of the bins were set to be evenly spaced between the maximum and minimum values of the covariate [ho2018package].
@@ 2.3.5 Propensity score matching
For a comprehensive introduction of matching-based methods, including PSM, refer to the overview by Stuart [stuart2010matching].
@@ 2.3.8 Targeted maximum likelihood estimator
The application of TMLE for estimating ATT involves the following steps [van2011targeted, fang2020two]:
@@
Apply the targeted learning theory to update the initial estimates of and to and respectively. These updates are performed such that they solve the efficient score equation:
where is defined as
and is the empirical estimator of .
@@
The TMLE estimator for the ATT is then obtained as:
In this study, we implemented TMLE by using the SuperLearner for the estimation of both propensity scores and outcomes [gruber2006package]. The model is denoted as TMLE_SL. In addition, we implemented a model, denoted as AIPW_SL, on the basis of the AIPW method and the SuperLearner’s estimation of the propensity scores and outcomes.
@@ 2.4 Simulation
In Scenarios 1 and 2, we assumed three independent covariates that followed the standard normal distribution. was a confounder that affected both the treatment selection and the outcome, and and affected the treatment selection and the outcome, respectively. In Scenario 3, an additional confounder was assumed to follow the standard normal distribution and to be independent of .
@@
Setting 3: Outcome model with the interaction term between the treatment variable and the covariate.
@@
Estimated the ATT by using the methods described in Section 2.3. For each method, we included only covariates for estimation. We used the logistic regression based on to estimate the PS for PSM, PSM_1:2, IPW, and AIPW. And we used the LR based on to estimate the outcomes for AIPW.
@@
The type I error rate, i.e., the average rate of rejecting the null hypothesis of versus the alternative hypothesis of using a two-sided test at the 0.05 significance level by making in the outcome models described in Table 2.
@@
For this study, all statistical analyses were conducted in R (Version 4.2.2) using the following libraries: MatchIt (Version 4.5.5), PSweight (Version 1.1.8), tmle (Version 2.0.0), and their respective dependencies.
@@ 3.1 Scenario 1: no violation of the identifiability assumptions
The distribution of the PS values for Scenario 1 assuming no violation of the positivity or the unconfoundedness is shown in Figure S1 in the Supplementary Material. As shown in the figure, the distributions are bounded away from 0 and 1, and thus the positivity assumption can be considered reasonable. The bias, empirical SD (hereafter referred to as SD), average theoretical SD, MSE, and type I error rate for the 10 methods in Scenario 1 under three different settings are shown in Figures S2-S4 in the Supplementary Material, respectively.
@@
In Setting 3 with the interaction between the treatment variable and the covariate in the true outcome model, most methods yield very small biases. In terms of SD, the matching-based methods in general have relatively larger SDs than the other methods. The MSE pattern is similar to the SD pattern. Under the null treatment effect, the outcome models for Settings 1 and 3 are the same, therefore, the findings regarding the type I error are also the same for these two settings.
@@ 3.2 Scenario 2: violation of the positivity assumption
The distribution of the PS values for Scenario 2 assuming violation of the positivity or the unconfoundedness is shown in Figure S1 in the Supplementary Material. As shown in the figure, a large proportion of subjects’ PS values are distributed at or close to 0 or 1 and thus the positivity assumption is considered violated. The bias, empirical SD, average theoretical SD, MSE, and type I error rates for the 10 methods in Scenario 2 under three different settings are shown in Figures S5-S7 in the Supplementary Material, respectively.
@@
In Setting 1 with linear effect in the true outcome models, CEM2 and IPW in general have noticeably larger biases compared to other methods. The increased bias of CEM2 in Scenario 2 compared to Scenario 1 is attributed to the impact of violating the positivity assumption. Under Scenario 2, many subjects cannot be matched exactly but are still matched when using two bins, thus causing a larger bias. The IPW also produces a relatively larger bias under Scenario 2. The SD pattern in Scenario 2 differs from that in Scenario 1 in the following ways: 1) The SDs of IPW, AIPW, and TMLE_SL have a noticeable increase, and 2) the average theoretical SD of TMLE_SL is much smaller than the empirical SD. In terms of MSE, LR and AIPW_SL are the best two performing models. Under the null treatment effect, CEM2, IPW, and TMLE_SL experience remarkable type I error inflation. According to Figure S5 in the Supplementary Material, the type I error inflation of the CEM2 or IPW is mainly due to the large bias, while the type I error inflation of the TMLE_SL is mainly due to its underestimation of the SD. In contrast, PSM and PSM_1:2 perform best in terms of type I error control.
@@
In Setting 2 with nonlinear effect in the true outcome models, in addition to the CEM2 and IPW, which are already substantially biased in Setting 1, LR and AIPW are also largely biased. The bias of CEM5 is relatively large when the prevalence rate of treatment is low and relatively small when the prevalence rate of treatment is high. The biases of PSM and PSM_1:2 are the smallest. The SDs of most models in Setting 2 increase compared to Setting 1, especially for IPW and LR. The AIPW_SL has a relatively small bias and the smallest SD, resulting the smallest MSE. Under the null treatment effect, LR, CEM2, IPW, AIPW, and CEM5 show severe type I error inflation, mainly due to their large biases (Figure S6 in the Supplementary Material); TMLE_SL also shows severe type I error inflation, mainly due to its underestimation of the SD (Figure S6 in the Supplementary Material). AIPW_SL has relatively mild type I error inflation, while PSM, PSM_1:2, and MDM have the best type I error control.
@@
In Setting 3 with the interaction between the treatment variable and the covariate in the true outcome model, the IPW, LR, and matching-based methods in general have relatively large biases (Figure S7 in the Supplementary Material). The doubly robust methods (i.e., AIPW, AIPW_SL, and TMLE_SL) in general have relatively small biases. Among the doubly robust methods, AIPW and TMLE_SL have relatively large SDs and hence relatively large MSEs. The MSE of AIPW_SL is the smallest, due to its relatively small bias and SD. The findings for Setting 3 regarding the type I error inflation are the same as those for Setting 1 (Figure S7 in the Supplementary Material).
@@ 3.3 Scenario 3: violation of the unconfoundedness assumption
For all methods and three settings of the outcome model, violation of the unconfoundedness assumption in Scenario 3 leads to a certain amount of bias (or a larger bias if the method is already biased) compared to Scenario 1 where this assumption is not violated (Figures S8-S10 in the Supplementary Material). The biases induced by violation of unconfoundedness assumption are all positive under our simulation settings. We conducted post-hoc analyses by varying the coefficient values of the unmeasured confounder in the three settings and found that the sign and magnitude of the induced bias are mainly determined by the coefficient value of the unmeasured confounder.
The pattern of SD is similar between Scenario 1 and Scenario 3. The MSEs observed in Scenario 3 are greater than those observed in Scenario 1. Under the null hypothesis, the type I error rates in Scenario 3 generally increase compared to Scenario 1 due to the larger biases.
@@ 4.1.1 Matching-Based Methods
Among matching-based methods, our simulation results show that the PSM model in general results in a small bias when the positivity assumption is held. When the positivity assumption is violated and when there is interaction between the treatment variable and the covariate (i.e., Scenario 2 Setting 3), a mild bias can be induced. This is because under the violation of the positivity assumption, the integration of in equation (2) is carried out in the overlapped range of the PS distributions for the two groups and hence the causal effect of interest has been deviated from the ATT. Similar issues occur in the other matching-based methods that we studied. Of note, the bias of the PSM can still be small in cases with positivity violation but without interaction between the treatment variable and the covariate, e.g., Scenario 2 Settings 1 or 2, because in equation (2), is a constant, and its integration over is the same as its integration in the overlapped range of the PS distribution.
@@
When considering the choice of matching ratio, our simulation indicates that 1 to 2 matching provides a reasonable balance between bias reduction and precision for estimating the ATT in most of our settings with the prevalence rate of treatment . This finding aligns with the simulation results of Austin [austin2010statistical] for observational studies, which demonstrated that matching one or two untreated subjects to each treated subject generally provides good performance for ATT estimation. For scenarios where the control group is substantially larger than the treatment group and high-quality matches are available, using 1 to r matching with r 2 may be considered to improve efficiency, though maintaining match quality is needed.